OpenStudy (anonymous):

Team X and Team Y have a tug of war. From their starting positions Team X pulls Team Y forward 3 meters, and Team X is then pulled forward 5 meters. Team Y then pulls Team X forward 2 meters. If the first team to be pulled forward 10 meters loses, how many more meters must Team Y pull Team X forward to win?

OpenStudy (anonymous):

@Disco619

OpenStudy (anonymous):

hold up let me read. lol

OpenStudy (disco619):

WAIT WAIT SORY SOOORY! I made a little MISTAKE! WAAAAAIT

OpenStudy (disco619):

LET ME MAKE THE DIAGRAM AGAIN!

OpenStudy (anonymous):

ok ok calm down.. you dont have to make the diagram you can explain in words if you want :)

OpenStudy (disco619):

Diagrams are easy... w8

OpenStudy (disco619):



OpenStudy (anonymous):

so how many does Y need to pull to get ahead?

OpenStudy (disco619):

So, In TOTAL. Team X is pulled forward: 5 - 3 + 2 = 4 meters. Total meters needed to pull Team X to win = 10 Meters needed = 10 - 4 = 6 meters [ANS]

OpenStudy (disco619):

Team Y needs to pull Team X 6 meters for them to win.

OpenStudy (anonymous):

Thank You!

OpenStudy (disco619):

Anytime... :)